# StreetLens: Enabling Human-Centered AI Agents for Neighborhood Assessment from Street View Imagery

### v1

@@ Abstract.
Traditionally, neighborhood studies have employed interviews, surveys, and manual image annotation guided by detailed protocols to identify environmental characteristics, including physical disorder, decay, street safety, and sociocultural symbols, and to examine their impact on developmental and health outcomes. While these methods yield rich insights, they are time-consuming and require intensive expert intervention. Recent technological advances, including vision-language models (VLMs), have begun to automate parts of this process; however, existing efforts are often ad hoc and lack adaptability across research designs and geographic contexts. In this demo paper, we present StreetLens, a human-centered, researcher-configurable workflow that embeds relevant social science expertise in a VLM for scalable neighborhood environmental assessments. StreetLens mimics the process of trained human coders by grounding the analysis in questions derived from established interview protocols, retrieving relevant street view imagery (SVI), and generating a wide spectrum of semantic annotations from objective features (e.g., the number of cars) to subjective perceptions (e.g., the sense of disorder in an image). By enabling researchers to define the VLM’s role through domain-informed prompting, StreetLens places domain knowledge at the core of the analysis process. It also supports the integration of prior survey data to enhance robustness and expand the range of characteristics assessed across diverse settings. We provide a Google Colab notebook to make StreetLens accessible and extensible for researchers working with public or custom SVI datasets. StreetLens represents a shift toward flexible, agentic AI systems that work closely with researchers to accelerate and scale neighborhood studies.
@@
journalyear: 2025copyright: acmlicensedconference: The 33rd ACM International Conference on Advances in Geographic Information Systems; November 3–6, 2025; Minneapolis, MN, USAbooktitle: The 33rd ACM International Conference on Advances in Geographic Information Systems (SIGSPATIAL ’25), November 3–6, 2025, Minneapolis, MN, USAprice: 15.00doi: -isbn: -$c$s: Human-centered computing Interactive systems and tools
@@ 1. Introduction
Figure 1. System workflow with StreetLens architecture. StreetLens consists of four modules (abbreviated as M#) and input data sources (abbreviated as I#).
This is the author’s version of the work.
@@
Neighborhood environments influence human well-being and have become a focus of research in urban studies, public health, and family social science (Pasco and White, 2024). Assessing neighborhood environmental features is important for understanding how neighborhood contexts affect outcomes like adolescent development (Pasco and White, 2024), mental health (Liu et al., 2020), and social cohesion (Bjornstrom and Ralston, 2014; Liu et al., 2020). Neighborhood environment assessments have relied on qualitative and semi-quantitative methods, including surveys (e.g., Los Angeles Family and Neighborhood Survey) (Bjornstrom and Ralston, 2014) and interviews (Pasco and White, 2024). As street view imagery (SVI) becomes more accessible, researchers have created systematic social observation (SSO) protocols (Odgers et al., 2012b; Pasco and White, 2020) to evaluate neighborhood environments using structured visual analysis using SVI. Protocols involve multiple human coders applying detailed codebooks to identify visual cues, such as signs of physical disorder, physical decay, and street safety. In practice, human coders rarely produce identical annotations for the same image, so multiple coders are needed per image, along with statistical analysis to compare their results. Despite offering invaluable insights, the process remains labor-intensive, dependent on trained experts, and difficult to scale across a wide variety of study contexts or geographic regions.
@@
Recent advancements in vision-language models (VLMs) have enabled researchers to automate certain aspects of the neighborhood environmental assessment process. Although researchers have begun to use VLMs with SVI (Liang et al., 2024; Huang et al., 2024), these efforts often apply models to a single image annotation task, such as identifying objects without a structured framework for adapting the models to comprehensively assess neighborhood environments. A limitation is the lack of a systematic method to “teach” VLMs in the way human coders are trained, including learning from the literature, coding protocols, and annotated examples. In addition, VLM results are often accepted as-is, without providing feedback to the researcher.
@@
To address this gap, we introduce StreetLens, an end-to-end researcher-centered workflow that mimics general human coder training processes. StreetLens guides VLM through reviewing relevant studies, studying coding manuals, examining example annotations, and comparing model output with those of experienced coders. StreetLens automates the assessment pipeline using open-source VLMs and supports flexible adaptation to different study contexts and geographic regions. Specifically, StreetLens enables researchers to embed domain knowledge as a central component of the workflow by leveraging relevant studies to configure the VLM through role prompting (Schulhoff et al., 2024), explicitly defining the VLM’s role in the assessment process. StreetLens then processes questions from the codebook (e.g., assessing physical disorder or identifying objects), retrieves relevant SVI, and generates structured annotations aligned with expert-defined protocols. StreetLens supports a feedback loop that allows researchers to compare and interpret automated coding with human coding, further facilitating the refinement of the research process. To support accessibility, we provide a Google Colab notebook that enables users to run StreetLens with either publicly available or user-provided image data, eliminating the need for advanced technical expertise. The result is a flexible and reusable workflow that facilitates assessing neighborhood environmental features and benefits various studies across geographic settings.
@@ 2. Case Study
This section introduces a case study from family social science that motivates the workflow of StreetLens. The case study (Pasco and White, 2020) aims to assess how neighborhood environments relate to adolescents’ ethnic and racial label usage. To match labels from semi-structured interviews with potential environmental features in neighborhoods, Pasco and White (2020) leverages the established SSO protocol (Odgers et al., 2012b) (i.e., a detailed set of instructions) to train human coders. Multiple trained coders then use Google Earth Pro to virtually walk through each street segment, evaluating the degree of physical decay (e.g., deteriorated buildings, poor sidewalks) and sociocultural symbols (e.g., Spanish-language signs, Latino-owned businesses).
After coding environmental features, Pasco and White (2020) assess the reliability of ratings by comparing coders’ assessments of the street segments using intraclass correlation coefficients. Such a validation step removes outliers and ensures consistent ratings across other coders while minimizing personal bias. Involving multiple coders accounts for variations in individual perception and judgment, thus enhancing the overall quality of the assessments. Therefore, StreetLens aims to serve as an additional agentic coder by enabling researchers to create an automated coder with domain-specific materials.
@@ 3. StreetLens
StreetLens provides a researcher-oriented workflow (Figure 1), targeting users whose research involves neighborhood environmental assessments. A researcher begins by using StreetLens through a simple, guided interface. The first module, M1. Data Processor, prompts the researcher to upload key materials such as codebooks, protocols, related papers, and example annotations, as well as to specify the study area for retrieving SVI data. This module organizes and prepares all inputs for the next steps. Next, M2. Automated Prompt Tuning uses the collected domain knowledge to define the role of the VLM agent and generate protocol-aligned prompts that follow the researcher’s coding instructions. These prompts are passed to M3. VLM Processor, which analyzes street-level images and generates the assessments of environmental features. The researcher then reviews the results with the examples created by human coders. Finally, M4. Feedback Provider allows the VLM agent to communicate back to the researcher, offering explanations on how the VLM agent interpreted the coding instructions. This feedback helps the researcher understand the agent’s reasoning.
@@
StreetLens starts by asking the researcher to choose a study area, such as a city or specific census tracts, where they want to assess environmental features. Based on the selected area, the system returns a set of predefined point locations. These points are derived from U.S. Census TIGER road data, which has been segmented at 5-meter intervals to support efficient data retrieval. For this demonstration, StreetLens uses Google Street View imagery to match the source used in the original case study with human annotations. Once the study area is set, the system prompts the researcher to indicate which materials are available, such as a codebook, protocol, sample annotations, or academic papers in the domain that the user wants StreetLens to focus.
@@
StreetLens generates a role by reviewing related papers and assigns a role to the VLM to help it “think” like a trained human coder. This step enhances the VLM’s performance (Schulhoff et al., 2024) and aligns with how human coders learn by reading background studies and understanding what to look for in the environment. To achieve this, StreetLens uses the following prompt with the abstracts of relevant papers to the large language model (LLM):
@@
You are an expert in the following fields and the author of the paper abstracts provided here: [I3. Abstracts of related papers]. Based on the expertise demonstrated, generate a general professional role description of yourself in one to two sentences, starting with ”You are” written in the second person. This will be used as a system prompt introduction.
For example, using two studies on neighborhood environments and Mexican-origin adolescents (Pasco and White, 2020, 2024), the LLM generates:
You are an expert in conducting mixed-methods research in urban sociology and ethnic studies, focusing on the impact of neighborhood environments on social behaviors and identity formation among Latinx adolescents. You specialize in using systematic social observations and qualitative interviews to compare and contrast the perspectives of researchers and adolescents, providing critical insights into the shared and unique aspects of urban environments and ethnic-racial identity within ethnically/racially segregated neighborhoods.
After defining the specific role for the VLM agent, StreetLens processes the codebook collected in M1, which contains questions and answer options for each environmental feature. For each feature, StreetLens uses the following prompt:
[I1. Question and answer options from codebook] Review the question and answer options above. Guide a vision-language model to assess environmental features in S using only visual input. First, write one sentence to complete the system prompt. Then, write 2–3 clear sentences for the user prompt using the EXACT SAME numeric options.
Finally, StreetLens asks if the user wants to modify the prompt using the materials collected in M1. For example, for the case study, protocols include detailed instructions for subjective assessments. We demonstrate how researchers can use the corresponding image-answer pairs to improve the VLM agent’s performance by applying in-context learning techniques. Similarly, researchers can incorporate existing annotations from human coders as additional image-answer pairs to improve the model. The following is an example of a generated prompt for assessing the condition of a sidewalk (corresponding to the “Decay 2” code theme in the SSO protocol (Pasco and White, 2020)):
[Assigned role prompt] Assess the environmental condition of the sidewalk in the street view image provided, focusing on its surface and overall maintenance. Evaluate the condition of the sidewalk based on visual cues from the image and choose the best matching option. Options are: 1 - Good (NO holes, sizable cracks, or crumbling or uneven pavement) 2 - Fair (Holes, sizable cracks, or crumbling or uneven pavement, outgrown weeds along SOME of the side walk) 3 - Poor (Holes, sizable cracks, or crumbling or uneven pavement, outgrown weeds along most or ALL of the sidewalk) 99 - Under construction or cannot Evaluate. Your response must be an integer. DO NOT PROVIDE ANY OTHER OUTPUT TEXT OR EXPLANATION.
Figure 2 shows an example prompt generated by StreetLens based on input from the codebook, protocol, and relevant academic papers.
@@
With the generated prompts from M2 and the retrieved SVI from M1, StreetLens generates assessments for each environmental feature. In this demo, StreetLens loads the open-source VLM InternVL32B (Zhu et al., 2025), which combines the image encoder InternViT-300M-448px-V2_5 with the language model Qwen2.5-1.5B. As the prompt includes SVI as well, StreetLens generates the ¡image¿ tags matching the number of processed SVI per prompt.
@@
After the VLM agent assesses the environmental features, StreetLens provides feedback to researchers that includes explanations of the agent’s assessments, leveraging reasoning capabilities. Moreover, the agent’s output acts as an additional coder, enabling the calculation of inter-rater reliability metrics. In the demo, StreetLens computes the intraclass correlation coefficient to validate the agreement between coders.
Figure 2. An example of StreetLens ’s automated prompt tuning procedure demonstrates how the system utilizes related academic papers, codebooks, and assessment protocols to generate domain-specific prompt.
@@ 4. Related Work
Assessing neighborhood environment characteristics has traditionally involved trained coders conducting SSO and in-person audits. For example, a seminal study by Sampson and Raudenbush (1999) employed video-based ratings of over 23,000 street segments in Chicago to quantify physical disorder. To scale such efforts, researchers later adopted “virtual audits” using platforms like Google Earth and Street View (Clarke et al., 2010; Rundle et al., 2011; Odgers et al., 2012a). These methods enabled remote assessment of neighborhoods with high inter-rater reliability and reduced cost. Building on this, computer vision techniques have been used to detect and quantify physical features such as urban greenery and sidewalk quality (Biljecki and Ito, 2021) and to infer higher-level, conceptual attributes like perceived safety from imagery (Naik et al., 2014; Biljecki and Ito, 2021). However, such supervised approaches are often constrained by task-specific labels and limited generalizability to new contexts. Recent advances in VLMs offer more flexible solutions by enabling open-ended scene interpretation through joint image-text representations. VLM-based methods have been used to assess walkability (Blečić et al., 2024) and generate structured descriptions of urban environments (Perez and Fusco, 2025), showing promise in automating scalable, semantically rich neighborhood audits across diverse settings. In contrast to prior work tailored to specific studies, StreetLens is a VLM-based, end-to-end, researcher-centered SSO workflow that simulates the human coder training process, enhancing adaptability across diverse study designs and geographic contexts.

### v2

@@ Abstract.
Traditionally, neighborhood studies have used interviews, surveys, and manual image annotation guided by detailed protocols to identify environmental characteristics, including physical disorder, decay, street safety, and sociocultural symbols, and to examine their impact on developmental and health outcomes. Although these methods yield rich insights, they are time-consuming and require intensive expert intervention. Recent technological advances, including vision language models (VLMs), have begun to automate parts of this process; however, existing efforts are often ad hoc and lack adaptability across research designs and geographic contexts. In this paper, we present StreetLens, a user-configurable human-centered workflow that integrates relevant social science expertise into a VLM for scalable neighborhood environmental assessments. StreetLens mimics the process of trained human coders by focusing the analysis on questions derived from established interview protocols, retrieving relevant street view imagery (SVI), and generating a wide spectrum of semantic annotations from objective features (e.g., the number of cars) to subjective perceptions (e.g., the sense of disorder in an image). By enabling researchers to define the VLM’s role through domain-informed prompting, StreetLens places domain knowledge at the core of the analysis process. It also supports the integration of prior survey data to enhance robustness and expand the range of characteristics assessed in diverse settings. StreetLens represents a shift toward flexible and agentic AI systems that work closely with researchers to accelerate and scale neighborhood studies. StreetLens is publicly available at https://knowledge-computing.github.io/projects/streetlens.
@@
$c$s: Human-centered computing Interactive systems and toolsfootnotetext: © Jina Kim et al. 2025. This is the author’s version of the work. It is posted here for personal use. Not for redistribution. The definitive version was published in The 1st ACM SIGSPATIAL International Workshop on Human-Centered Geospatial Computing (GeoHCC ’25), https://doi.org/10.1145/3764917.3771334.
@@ 1. Introduction
Figure 1. Example of input data for StreetLens (top) and system workflow (bottom) illustrating StreetLens architecture. StreetLens comprises four modules (M#). Input and output data are denoted as I# and O#, respectively.
@@
Neighborhood environments influence human well-being and have become a focus of research in urban studies (Sampson and Raudenbush, 1999), public health (Bond et al., 2012), and family social science (Pasco and White, 2024). Assessing neighborhood environmental characteristics is crucial to understand how neighborhood contexts influence outcomes such as adolescent development (Pasco and White, 2024), mental health (Liu et al., 2020), and social cohesion (Bjornstrom and Ralston, 2014; Liu et al., 2020). Neighborhood environment assessments have relied on qualitative and semi-quantitative methods, including surveys (e.g., Los Angeles Family and Neighborhood Survey) (Bjornstrom and Ralston, 2014) and interviews (Pasco and White, 2024). As street view imagery (SVI) becomes more accessible, researchers have created systematic social observation (SSO) protocols (Odgers et al., 2012b; Pasco and White, 2020) to evaluate neighborhood environments using structured visual analysis using SVI. Protocols involve multiple human coders applying detailed codebooks to identify visual cues, such as signs of physical disorder, physical decay, and street safety. In practice, human coders rarely produce identical annotations for the same image, so multiple coders are needed per image, along with statistical analysis to compare their results. Despite offering invaluable insights, the process remains labor-intensive, dependent on trained experts, and challenging to scale across a wide variety of study contexts or geographic regions.
@@
Recent advances in vision language models (VLMs) have enabled researchers to automate certain aspects of the neighborhood environmental assessment process. Although researchers have begun to use VLMs with SVI (Liang et al., 2024; Huang et al., 2024; Jang and Kim, 2025), these efforts often apply models to simple image annotation tasks, such as identifying objects in the physical environment, without a structured framework for adapting the models to fully assess the characteristics of a neighborhood. A limitation is the lack of a systematic method to “teach” VLMs in the way human coders are trained, including learning from the literature, coding protocols, and annotated examples. Also, VLM results are often accepted as is, without providing feedback to the researcher.
@@
To address this gap, we introduce StreetLens, an end-to-end researcher-centered workflow that mimics general human coder training processes. StreetLens guides VLM through reviewing relevant studies, studying coding manuals, examining example annotations, and comparing the model output with those of experienced coders. StreetLens automates the assessment pipeline using open-source VLMs and supports flexible adaptation to different study contexts and geographic regions. Specifically, StreetLens allows researchers to incorporate domain knowledge as a central component of the workflow by leveraging relevant studies to configure the VLM through role prompting (Schulhoff et al., 2025), explicitly defining the role of the VLM in the assessment process. StreetLens then processes the codebook questions (for example, assessing physical disorder or identifying objects), retrieves the relevant SVI, and generates structured annotations aligned with expert-defined protocols. To support accessibility, we provide a Google Colab notebook that allows users to run StreetLens with publicly available or user-provided image data, eliminating the need for advanced technical expertise. The result is a flexible and reusable workflow that facilitates assessing neighborhood environmental characteristics and benefits various studies across geographic settings.
@@ 2. Case Study
This section introduces a case study from family social science that motivates the workflow of StreetLens. The case study (Pasco and White, 2020) aims to assess the relationship between neighborhood environments and the use of ethnic and racial labels (i.e., defining social identity) by adolescents. To match ethnic and racial labels from semi-structured interviews with potential environmental features in neighborhoods, Pasco and White (2020) leverage the established SSO protocol (Odgers et al., 2012b) (i.e., a detailed set of instructions) to train human coders. Multiple trained coders then use Google Earth Pro to virtually walk through each street segment, evaluating the degree of physical decay (e.g., deteriorated buildings, poor sidewalks) and sociocultural symbols (e.g., Spanish-language signs, Latino-owned businesses).
After coding environmental features, Pasco and White (2020) assess the reliability of the ratings by comparing coders’ assessments of the street segments using intraclass correlation coefficients. This validation step removes outliers and ensures consistent ratings among coders, while minimizing personal bias. Involving multiple coders accounts for variations in individual perception and judgment, thus enhancing the overall quality of the assessments. Therefore, StreetLens aims to serve as an additional agentic coder by enabling researchers to create an automated coder with domain-specific materials.
@@ 3. StreetLens
StreetLens provides a researcher-oriented workflow (Figure 1), targeting users whose research involves neighborhood environmental assessments. A researcher begins using StreetLens through a simple and guided interface. The first module, M1. Data Processor, prompts the researcher to upload key materials such as codebooks, protocols, related publications, and example annotations, as well as to specify the study area to retrieve the SVI data. This module organizes and prepares all inputs for the next steps. Next, M2. Automated Prompt Tuning uses the collected domain knowledge to define the role of the VLM agent and generate protocol-aligned prompts that follow the researcher’s coding instructions. These prompts are passed to M3. VLM Processor, which analyzes street-level images and generates assessments of environmental characteristics. The researcher then reviews the results using previous coding created by human coders. Finally, M4. Feedback Provider allows the VLM agent to provide explanations on how the VLM agent interpreted the coding instructions. This feedback helps the researcher understand the reasoning of the agent.
@@
StreetLens starts by asking the researcher to choose a study area, such as a city or specific census tracts, where they want to assess environmental characteristics. Based on the selected area, the system returns a set of predefined point locations. These points are derived from U.S. Census TIGER road data, which have been sampled at 5-meter intervals to support efficient data retrieval. For this paper, StreetLens uses Google Street View imagery to match the source used in the original case study (Section 2) with human annotations. Once the study area is set, the system prompts the researcher to indicate which materials are available, such as a codebook, protocol, sample annotations, or academic papers in the domain that the user wants StreetLens to focus on.
@@
StreetLens generates a role by reviewing related papers and assigns a role to the VLM to help it “think” like a trained human coder. This step enhances the VLM’s performance (Schulhoff et al., 2025) and aligns with how human coders learn by reading background studies and understanding what to look for in the environment. To achieve this, StreetLens uses the following prompt with the abstracts of relevant papers to the large language model (LLM):
@@
You are an expert in the following fields and the author of the paper abstracts provided here: [I2. Abstracts of related papers]. Based on the expertise demonstrated, generate a general professional role description of yourself in one to two sentences, starting with ”You are” written in the second person. This will be used as a system prompt introduction.
After defining the specific role for the VLM agent, StreetLens processes the codebook collected in M1, which contains questions and answer options for each environmental feature. StreetLens first determines whether each pair of questions and answers presents a subjective or objective task. This step allows StreetLens to align the VLM agent with the relevant task category. For example, in this case study (Section 2), StreetLens distinguishes whether a given question and answer pair aims to assess subjective neighborhood perceptions or to verify the presence of specific objects (e.g., graffiti or bike lanes). StreetLens uses the following prompt for this procedure:
You are a classifier of annotation tasks. Given a question and its answer options, decide if the task is perception (holistic/qualitative scene judgment such as condition/quality/intensity ratings) or object_detection (presence, counting, or localization of specific object instances). Rules: If it asks to rate/assess overall condition or quality (e.g., Good/Fair/Poor), label as perception. If it asks to detect, count, or verify specific objects (e.g., cars, signs, pedestrians), label as object_detection. [I4. Question and answer options from codebook] Return only a single integer: 0 if perception, 1 if object_detection. Do not include any words, JSON, spaces, or punctuation.
Then, for each question and its corresponding answer options, StreetLens applies the following prompt to generate a codebook prompt:
Instruction: Rewrite the question as a clear, self-contained sentence, prefixed with ”Question:”. Then, rewrite each answer option as a full sentence explaining the meaning, starting with its number. Keep all numbers and meaning intact. Output plain text only, one sentence per line. [I4. Question and answer options from codebook]
For example, Figure 2 shows the generated O2. Role Prompt and O3. Codebook Prompt. In this case, the LLM takes as input two studies on neighborhood environments and Mexican-origin adolescents (Pasco and White, 2020, 2024), along with the ‘Disorder 3’ question and its answer options from the codebook.
Figure 2. StreetLens’s automated prompt-tuning output (a domain-specific prompt) using related papers and the codebook.
Figure 3. StreetLens ’s M3. Vision Language Model Processor and M4. Feedback Provider, showing both the input and the processed output. Refer Figure 2 for O2. Role Prompt.
@@
With the M2-generated prompts and the SVIs collected from M1, StreetLens generates assessments for each environmental feature. The current version of StreetLens loads the open-source lightweight VLM InternVL3-2B (Zhu et al., 2025), which consists of an image encoder InternViT-300M-448px-V2_5 with a language model Qwen2.5-1.5B. Figure 3 presents how StreetLens encodes a set of street view images using the image encoder. After the projection layer, StreetLens passes the resulting image token embeddings, together with the tokenized text inputs from M2 prompts (i.e., O2 and O3), to the language model. In addition to the prompts generated in M2, StreetLens leverages image–answer pairs from the I3 protocol, used as training materials for human coders, to perform in-context learning.
@@
After the VLM agent evaluates the environmental characteristics, StreetLens provides feedback to researchers that includes explanations of the agent’s evaluations, utilizing reasoning capabilities. For example, when coding the ‘Decay 1’ measure (see Figure 1 for measure details), which assesses the general condition of visible streets, StreetLens provides the explanation of a certain street block ID #281 assessment, ‘There are only slight cracks, and any potholes present have been fixed or covered’ (Figure 3). This process improves the interpretability of StreetLens.
For demonstration, StreetLens is available at https://knowledge-computing.github.io/projects/streetlens, integrating data from the original case study (Section 2) and connecting to Cloudflare through a server located at the University of Minnesota. Users can individually toggle each module using buttons to explore StreetLens.
@@ 4. Related Work
Assessing neighborhood environment characteristics has traditionally involved trained coders conducting SSO and in-person audits. For example, a seminal study by Sampson and Raudenbush (1999) employed video-based ratings of more than 23,000 street segments in Chicago to quantify physical disorder. To scale such efforts, researchers later adopted “virtual audits” using platforms like Google Earth and Street View (Clarke et al., 2010; Rundle et al., 2011; Odgers et al., 2012a). These methods enabled remote assessment of neighborhoods with high inter-rater reliability and reduced cost. Based on this, computer vision techniques have been used to detect and quantify physical characteristics such as urban greenery and sidewalk quality (Biljecki and Ito, 2021) and to infer conceptual attributes of a higher level, such as perceived safety from images (Naik et al., 2014; Biljecki and Ito, 2021). However, such supervised approaches are often constrained by task-specific labels and limited generalizability to new contexts. Recent advances in VLMs offer more flexible solutions by enabling open-ended scene interpretation through joint image-text representations. VLM-based methods have been used to assess walkability (Blečić et al., 2024), generate structured descriptions of urban environments (Perez and Fusco, 2025), and detect objects within audit categories (e.g., tree, vehicle, trash bin) (Jang and Kim, 2025), demonstrating promise in automating scalable, semantically rich neighborhood audits across diverse settings. In contrast to prior work tailored to specific studies, StreetLens is a VLM-based, end-to-end, researcher-centered SSO workflow that simulates the human coder training process.